For y = 4x2 * ex on the interval [-4,4]

a) find the absolute max/min :

b) find the extreme slopes

c) make the y’ and y’’ number lines and show their A/B patterns

d) List the local max and mins and IPs from c)

e) Sketch y showing all the important points and label its patterns, and list its xintercepts and y intercept.
Answer #1

Sel y=4x² * on interval [-4, 4] y = 4(x² et + ² x an] =0 ex ( 22 + 2x] so e zo Innehno a2+ 2n-o r (+2) a 220 p = 2 Criticalat x = 0 4 = 4xo e so at &- -2 4-4x+2)2. & 4x4x e - 2016.0 02 to pizo point is Absolute minima. y =o abs min Maximan value wiLocal minima =(-2, 2-165) Absolute minima = (oo) Absolute masline = (4 tends too) at Asyutostre(-2, 2.165) (0.0)(-2, 2.165) (0.0)